OpenStudy (rainbow_rocks03):

what is 0.05208333333 as a fraction because I have to divide 5/8 divided by 3/4 and I got 0.05208333333 as my answer but I need it as a fraction.

OpenStudy (rainbow_rocks03):

@DanJS

OpenStudy (kendricklamar2014):

\[\frac{ 5 }{ 8 } \div \frac{ 3 }{ 4 } = \frac{ 5 }{ 96 } (0.052083...)\] @rainbow_rocks03

OpenStudy (rainbow_rocks03):

how did you get that?

OpenStudy (kendricklamar2014):

by dividing

OpenStudy (rainbow_rocks03):

also that is not an answer choice

OpenStudy (kendricklamar2014):

What are the answer choices?

OpenStudy (rainbow_rocks03):

A. 15/32 B. 5/6 C. 1 1/5 D. 2 2/15

OpenStudy (rainbow_rocks03):

@DanJS

OpenStudy (mathstudent55):

To divide a fraction by a fraction, write the first fraction the way it is. Then change the division into multiplication. Then flip the second fraction. 

OpenStudy (mathstudent55):